Q1/ For the unit vectors â, and â shown below, find the scalars a and B such that ał, – 3, = Bſ. ܕܐܶ 60° 1 îi
Answer #1

Given, î= î and 2 = (costo)i + (Sinto) 13 2 2 Now, xâ-322 = ß a î - 3 - → 33 = ß If X = 32 -333 ß > 2 ñ ß = -353 2 x= 3/2 and